# Good Things Come in Pairs: Paired Autoencoders for Inverse Problems

### v1

@@ Abstract
In this book chapter, we discuss recent advances in data-driven approaches for inverse problems. In particular, we focus on the paired autoencoder framework, which has proven to be a powerful tool for solving inverse problems in scientific computing. The paired autoencoder framework is a novel approach that leverages the strengths of both data-driven and model-based methods by projecting both the data and the quantity of interest into a latent space and mapping these latent spaces to provide surrogate forward and inverse mappings. We illustrate the advantages of this approach through numerical experiments, including seismic imaging and classical inpainting: nonlinear and linear inverse problems, respectively. Although the paired autoencoder framework is likelihood-free, it generates multiple data- and model-based reconstruction metrics that help assess whether examples are in or out of distribution. In addition to direct model estimates from data, the paired autoencoder enables latent-space refinement to fit the observed data accurately. Numerical experiments show that this procedure, combined with the latent-space initial guess, is essential for high-quality estimates, even when data noise exceeds the training regime. We also introduce two novel variants that combine variational and paired autoencoder ideas, maintaining the original benefits while enabling sampling for uncertainty analysis.
@@ 1 Introduction
This book chapter discusses a novel class of data-driven methodologies for tackling inverse problems, leveraging the power of paired autoencoders. This framework centers on projecting both the observed data and the corresponding quantities of interest into their respective lower-dimensional latent spaces. By learning the intricate mappings between these latent representations, we effectively construct a surrogate model capable of approximating both the forward and inverse processes. This framework demonstrates significant performance gains compared to traditional data-driven approaches, yielding high-quality estimates for the recovered quantities of interest. In contrast to traditional data-driven machine learning approaches for inverse problems, paired autoencoders offer statistical measures to assess the reliability of the inverse solution, therefore addressing a key limitation of detecting out-of-distribution data. Moreover, variational-based paired autoencoder approaches extend this capability by providing uncertainty estimates for the inverse solution, as demonstrated through various numerical experiments.
@@
In Section 3, we explore extensions of the paired autoencoder framework aimed at incorporating uncertainty estimation using a variational paired autoencoder. Specifically, we examine two methodologies, including the use of a variational autoencoder as outlined in Section 3.1, its corresponding inference mechanism discussed in Section 3.1.1, and the integration of probabilistic modeling in the latent space via a variational encoder-decoder, as presented in Section 3.2.
To underscore the practical utility and advantages of paired autoencoders, we showcase their application to two distinct prototype problems. The first application focuses on simple image restoration tasks in Section 4.1, demonstrating the framework’s ability to recover clean images from noisy or degraded observations. The second application tackles a more complex problem in seismic inference in Section 4.3, highlighting the potential of paired autoencoders for extracting valuable subsurface information from seismic data. The network architectures and implementation details are presented in Appendix A. Finally, we conclude this work with a comprehensive discussion of our findings and outline promising future research directions within the burgeoning field of paired autoencoders in Section 5.
@@ 1.1 Inverse Problems
A key challenge in inverse problems is their ill-posedness, meaning solutions may not exist (no model perfectly explains the data), may not be unique (multiple models fit the data equally well), or may be unstable (small data perturbations cause large changes in the estimated model), [56]. Consequently, specialized techniques are essential. These include regularization, which introduces constraints or prior information to stabilize the solution; optimization, which frames the problem as minimizing a loss between observations forward model; and statistical methods, which incorporate statistical models of data and unknowns to quantify uncertainty [18]. The study of inverse problems is a vibrant and interdisciplinary field, drawing upon mathematics, physics, statistics, and computer science. It plays a crucial role in advancing our understanding of complex systems and enabling new technologies across various domains. Limitations of classical inversion techniques are evident in applications, where large observational errors and simplified mathematical models can lead to local minima or constitute multi-modal posterior distributions. Consequently, classical inversion techniques become impractical for such applications without additional processing. Advanced techniques are necessary to overcome these challenges.
@@ 1.2 Machine Learning Approaches for Inverse Problems
The emergence of data-driven approaches, particularly deep neural networks (DNNs), has significantly impacted the field of inverse problems [63, 29, 6, 55]. Traditionally, solving inverse problems relied heavily on analytical methods, numerical simulations, optimization algorithms, and sampling methods. Data-driven approaches aim to circumvent the challenges by utilizing those inversion methodologies. Below, we provide a brief overview of some deep learning approaches for inverse problems, to highlight some connections and differences to paired autoencoders.
@@
For instance, consider learning the mapping between measurements $y$ and model parameters $x$ directly from training data. By assuming a true underlying mapping $\Psi\colon\mathcal{Y}\to\mathcal{X}$, training a neural network on a large dataset of paired input-target samples (i.e., pairs of measurements $y$ and corresponding QoI $x$) may learn to approximate the inverse mapping, effectively bypassing the need for explicit inversion algorithms. To be more precise, these techniques train neural networks, such as $\Phi:\mathcal{Y}\times\mathcal{T}\to\mathcal{X}$, to directly map observed data $y$ onto estimates $\hat{x}=\Phi(y;\theta)$, where $\theta\in\mathcal{T}$ are trainable network parameters. To improve readability, we often omit neural networks’ explicit network parameter dependence, for instance, denoting them simply as $\Phi(\,\cdot\,)$. Given sufficient and representative training pairs $(x_{i},y_{i})_{i=1}^{N}$ for the input and target distributions $\pi_{y}$ and $\pi_{x}$, these end-to-end approaches have shown impressive results in producing reasonable reconstructions without explicit reliance on physical models. Because there is no unique, one-to-one relation between $x$ and $y$ for many inverse problems, the trained network implicitly contains implicit regularization: Given an input $y$, a non-stochastic network returns a single $\hat{x}$ out of all $x$ that generate $y$ under the forward operator $F$.
@@
$\min_{\theta}\ \mathbb{E}_{x\sim\pi_{x},y\sim\pi_{y}}\ \|x-\hat{x}_{k_{\text{max}}}(y;\theta)\|^{2}_{2},$ (3)
where $\hat{x}_{k_{\text{max}}}(y;\theta)$ is the output of the unrolled algorithm (2), which depends on the input data $y$ and learnable parameters. Each training iteration requires a pair $(x_{i},y_{i})_{j=1}^{J}$ of data and true QoI. Other versions of unrolled algorithms may (partially) learn the forward operator by unrolling the optimization algorithm as
@@
Upon the availability of larger training datasets and more powerful computational resources, we may aim to learn more from the data by training (deep) neural networks (DNN) with many parameters. Various approaches exist to learn a regularizer $\mathcal{R}$ as a deep neural network. In [88] the authors train the regularizer using examples images $x\sim\pi_{x}$ from the distribution of ground truth images, and images $x_{c}\sim\pi_{c}$ from the distribution of corrupted images. The idea is that a good regularization function assumes a large value for corrupted images, and a low value for correct images. Note that corrupted images are often generated by degrading good images, i.e., samples $x_{c}$ from $\pi_{c}$ are conditioned on $x$. The corresponding minimization problem reads
@@
$\min_{\theta}\mathbb{E}_{x_{c}\sim\pi_{c}|x\sim\pi_{x}}\|\Phi(x_{c};\theta)-x\|,$ (9)
@@
Assuming we have a forward differential operator, boundary conditions and spatial measurements $u(\chi_{j})=\tau_{j}$ for $j=1,\dots,m$, the solution of the PDE $u$ is modeled by a DNN, $\Phi(\,\cdot\,;\theta)$, parameterized by $\theta$. In the case of parameter estimation, these $\tau_{j}$ may coincide with our data $y_{i}$. Training simultaneously learns an approximation for the solution of the PDE $u(\chi)$ by $\Phi(\chi;\theta)$, and our QoI through minimization of the following quantities (if applicable): first, the mismatch computation of the forward model at the estimated solution; second, the violation of the predicted solution of boundary conditions; third the violation of the predicted solution of the physics governing the system; lastly, the mismatch between $u$ and $\tau_{j}$ evaluated at each $\chi_{j}$. Derivatives of $\Phi(\chi;\theta)$ can be computed using automatic differentiation, and mean-squared-error (MSE) is a common choice to compute these metrics.[28].
@@
Method $F$ for training $F$ for inference Data $y$ for training Paired $\{x,y\}$ for training
@@
Model-constrained approaches [95, 96] are similar to PINNs, but different in that they parametrize the inverse map by a DNN. Heuristically, these approaches learn the inverse mapping constrained by the forward map. Like PINNs, Model-constrained approaches attempt to ameliorate DNN data dependence by introducing additional information through regularization terms in optimization. Assuming a forward map $F$ is available, the model is aware that the observations it sees come from the forward model. Letting $\Phi(y;\theta)$ denote the learned inversion, model training optimizes
$\min_{\theta}\mathbb{E}_{x\sim\pi_{x},y\sim\pi_{y}}\left\|x-\Phi(y;\theta)\right\|_{2}^{2}+\alpha\left\|y-F(\Phi(y;\theta))\right\|_{2}^{2}$ (11)
@@
Several strategies exist to represent $\Phi$. For example, [96] introduces the use of encoder/decoder pairs, where the encoder learns the map from $x$ to observations $y$, and thus the decoder learns the inverse map.
@@ 1.3 Autoencoders
$x$$\begin{matrix}\mbox{encoder}\\ e\end{matrix}$$z$$\begin{matrix}\mbox{decoder}\\ d\end{matrix}$$x$
@@
The training of autoencoders is considered an unsupervised or self-supervised learning approach, where the network training learns to reconstruct the input data without relying on corresponding pre-defined labels or target data. Given input data $(x_{i})_{i=1}^{n_{x}}$, the autoencoder’s learning process involves minimizing a loss function between the original input and the reconstructed output. For instance, for the mean squared error (MSE) loss, we have
$\min_{\theta}\ \tfrac{1}{n_{x}}\sum_{i=1}^{n_{x}}\left\|d(e(x_{i};\theta^{e});\theta^{d})-x_{i}\right\|_{2}^{2}.$
@@ 1.4 Linear Autoencoder
In the case of linear autoencoders where both the encoder and decoder are linear transformations, i.e., $e=E\in\mathbb{R}^{r\times n}$ and $d=D\in\mathbb{R}^{n\times r}$ and each element of those matrices are trainable, with $r\leq n$ it is well known that the optimal solution is closely related to principal component analysis (PCA). This relationship has been established by assuming samples $x_{1},\ldots,x_{n_{x}}$ with $n_{x}\gg r$ and the corresponding data matrix $[x_{1},\ldots,x_{n_{x}}]$ has at least column rank $r$. The mathematical equivalence between PCA and linear autoencoders has been established in several works, notably [17, 9, 99]. These studies demonstrate that the weight matrices of the encoder and decoder correspond to the eigenvectors of the data covariance matrix associated with the largest eigenvalues. This insight has been leveraged in various applications, including dimensionality reduction, denoising, and feature extraction, where linear autoencoders provide an interpretable and computationally efficient framework for data representation. For further details, see [12], which explores regularized extensions of linear autoencoders, and [127], which discusses their use in model reduction.
A deeper understanding of optimal linear autoencoders may be gained by treating $x$ as a random variable [^4] To simplify notation, we will use lowercase $x$ to represent a random variable, relying on context to distinguish between random variables and samples., and analyzing the problem through the lens of the Bayes risk formulation, i.e.,
$\min_{\text{rank}(A)\leq r}\ \mathbb{E}\left\|Ax-x\right\|_{2}^{2}.$ (12)
Here, $\mathbb{E}$ denotes the expectation of a random variable $x$. Let us assume that $x$ has a finite second moment that is symmetric positive definite $\Gamma_{x}=L_{x}L_{x}^{\top}=\mathbb{E}\,xx^{\top}$ with a symmetric decomposition by a matrix $L_{x}\in\mathbb{R}^{n}$. In this case, the objective function in Equation 12 reduces to
$\mathbb{E}\left\|Ax-x\right\|_{2}^{2}$ $=\text{tr}\left((A-I)L_{x}L_{x}^{\top}(A-I)^{\top}\right)=\left\|AL_{x}-L_{x}\right\|_{\text{F}}^{2},$

### v2

@@ Abstract
In this work, we discuss recent advances in data-driven approaches for inverse problems. In particular, we focus on the paired autoencoder framework, which has proven to be a powerful tool for solving inverse problems in scientific computing. The paired autoencoder framework is a novel approach that leverages the strengths of both data-driven and model-based methods by projecting both the data and the quantity of interest into a latent space and mapping these latent spaces to provide surrogate forward and inverse mappings. We illustrate the advantages of this approach through numerical experiments, including seismic imaging and classical inpainting: nonlinear and linear inverse problems, respectively. Although the paired autoencoder framework is likelihood-free, it generates multiple data- and model-based reconstruction metrics that help assess whether examples are in or out-of-distribution. In addition to direct model estimates from data, the paired autoencoder enables latent-space refinement to fit the observed data accurately. Numerical experiments show that this procedure, combined with the latent-space initial guess, is essential for high-quality estimates, even when data noise exceeds the training regime. We also introduce two novel variants that combine variational and paired autoencoder ideas, maintaining the original benefits while enabling sampling for uncertainty analysis.
@@ 1 Introduction
This work discusses a novel class of data-driven methodologies for tackling inverse problems, leveraging the power of paired autoencoders. This framework centers on projecting both the observed data and the corresponding quantities of interest into their respective lower-dimensional latent spaces. By learning the intricate mappings between these latent representations, we effectively construct a surrogate model capable of approximating both the forward and inverse processes. This framework demonstrates significant performance gains compared to traditional data-driven approaches, yielding high-quality estimates for the recovered quantities of interest. In contrast to traditional data-driven machine learning approaches for inverse problems, paired autoencoders offer statistical measures to assess the reliability of the inverse solution, therefore addressing a key limitation of detecting out-of-distribution data. Moreover, variational-based paired autoencoder approaches extend this capability by providing uncertainty estimates for the inverse solution, as demonstrated through various numerical experiments.
@@
In Section 3, we explore extensions of the paired autoencoder framework aimed at incorporating uncertainty estimation using a variational paired autoencoder. Specifically, we examine two methodologies, including the use of a variational autoencoder as outlined in Section 3.1, its corresponding inference mechanism discussed in Section 3.2, and the integration of probabilistic modeling in the latent space via a variational encoder-decoder, as presented in Section 3.3.
To underscore the practical utility and advantages of paired autoencoders, we showcase their application to two distinct prototype problems. The first application focuses on simple image restoration tasks in Section 4.1, demonstrating the framework’s ability to recover clean images from noisy or degraded observations. The second application tackles a more complex problem in seismic inference in Section 4.3, highlighting the potential of paired autoencoders for extracting valuable subsurface information from seismic data. The network architectures and implementation details are presented in Appendix 6. Finally, we conclude this work with a comprehensive discussion of our findings and outline promising future research directions within the burgeoning field of paired autoencoders in Section 5.
@@ 1.1 Inverse Problems
A key challenge in inverse problems is their ill-posedness, meaning solutions may not exist (no model perfectly explains the data), may not be unique (multiple models fit the data equally well), or may be unstable (small data perturbations cause large changes in the estimated model), [63]. Consequently, specialized techniques are essential. These include regularization, which introduces constraints or prior information to stabilize the solution; optimization, which frames the problem as minimizing a loss between observations and forward modeled data; and statistical methods, which incorporate statistical models of data and unknowns to quantify uncertainty [21]. The study of inverse problems is a vibrant and interdisciplinary field, drawing upon mathematics, physics, statistics, and computer science. It plays a crucial role in advancing our understanding of complex systems and enabling new technologies across various domains. Limitations of classical inversion techniques are evident in applications, where large observational errors and simplified mathematical models can lead to local minima or constitute multi-modal posterior distributions. Consequently, classical inversion techniques become impractical for such applications without additional processing. Advanced techniques are necessary to overcome these challenges.
@@ 1.2 Machine Learning Approaches for Inverse Problems
The emergence of data-driven approaches, particularly deep neural networks (DNNs), has significantly impacted the field of inverse problems [71, 33, 8, 62, 104]. Traditionally, solving inverse problems relied heavily on analytical methods, numerical simulations, optimization algorithms, and sampling methods. Data-driven approaches aim to circumvent the challenges by utilizing those inversion methodologies. Below, we provide a brief overview of some deep learning approaches for inverse problems, to highlight some connections and differences to paired autoencoders. Furthermore, Table  1 provides a summary of applicability of the various approaches, based on the availability of a forward operator, observed data, and model-data pairs during training and inference.
@@
For instance, consider learning the mapping between measurements $y$ and model parameters $x$ directly from training data. By assuming a true underlying mapping $\Psi\colon\mathcal{Y}\to\mathcal{X}$, training a neural network on a large dataset of paired input-target samples (i.e., pairs of measurements $y$ and corresponding QoI $x$) may learn to approximate the inverse mapping, effectively bypassing the need for explicit inversion algorithms. To be more precise, these techniques train neural networks, such as $\Phi:\mathcal{Y}\times\mathcal{T}\to\mathcal{X}$, to directly map observed data $y$ onto estimates $\hat{x}=\Phi(y;\theta)$, where $\theta\in\mathcal{T}$ are trainable network parameters. To improve readability, we often omit neural networks’ explicit network parameter dependence, for instance, denoting them simply as $\Phi(\,\cdot\,)$. In the supervised learning scenario we sample pairs of measurements $y$ and corresponding QoI $x$ from the joint distribution $(x,y)\sim\pi$. Denote the input and target marginal distributions using $\pi_{y}$ and $\pi_{x}$. Given sufficient and representative training pairs $(x_{i},y_{i})_{i=1}^{N}$ from the joint distribution, these end-to-end approaches have shown impressive results in producing reasonable reconstructions without explicit reliance on physical models. Because there is no unique, one-to-one relation between $x$ and $y$ for many inverse problems, the trained network implicitly contains implicit regularization: Given an input $y$, a non-stochastic network returns a single $\hat{x}$ out of all $x$ that generate $y$ under the forward operator $F$.
@@
$\min_{\theta}\ \mathbb{E}_{(x,y)\sim\pi}\ \|x-\hat{x}_{k_{\text{max}}}(y;\theta)\|^{2}_{2},$ (3)
where $\hat{x}_{k_{\text{max}}}(y;\theta)$ is the output of the unrolled algorithm (2), which depends on the input data $y$ and learnable parameters. Each training iteration requires pairs $(x_{i},y_{i})_{j=1}^{J}$ of data and true QoI. Other versions of unrolled algorithms may (partially) learn the forward operator by unrolling the optimization algorithm as
@@
Upon the availability of larger training datasets and more powerful computational resources, we may aim to learn more from the data by training (deep) neural networks (DNN) with many parameters. Various approaches exist to learn a regularizer $\mathcal{R}$ as a deep neural network. In [100] the authors train the regularizer using (unpaired) examples images $x\sim\pi_{x}$ from the distribution of ground truth images, and images $x_{c}\sim\pi_{c}$ from the distribution of corrupted images. The idea is that a good regularization function assumes a large value for corrupted images, and a low value for correct images. Note that corrupted images are often generated by degrading good images, i.e., samples $x_{c}$ from $\pi_{c}$ are conditioned on $x$. The corresponding minimization problem reads
@@
An area of active research concerns Flow-Matching (FM) [3, 94] for generative modeling, where a learned space and time-dependent velocity field $v(x,t;\theta)$ maps a simple source distribution (often Gaussian) into the target data distribution. The likelihood of $x$ is computable using the trained flow [94], and may be used as the prior in Equation 6. The regularizer then reads [148]
$\Phi(x;\theta)_{\text{flow}}=\|x_{t=0}\|_{2}^{2}+\int_{t=0}^{t=1}\operatorname{tr}\bigg{(}\frac{\partial}{\partial x}v(x,t;\theta)\bigg{)}dt,$ (8)
which involves integrating an ODE back in time, starting at the current QoI estimate $x$, while computing the trace [56]. Here, the flow end-time $t=1$ corresponds to the target distribution, and $t=0$ corresponds to the source. The trace computation (approximation) and time-integration are typically considered computationally challenging [126, 41]; features that are shared by both diffusion models and flow matching.
@@
$\min_{\theta}\mathbb{E}_{(x_{c},x)\sim\pi}\|\Phi(x_{c};\theta)-x\|,$ (10)
@@
Learning Data-Dependent Flows, Conditional & Guided generation.
Flow-matching (FM, [3, 94]) recently gained popularity in the context of solving inverse problems. One of the reasons behind this is the various ways in which a flow trained via FM may be utilized. Here, we informally describe a few, while omitting some details. Some of the concepts in this paragraph also apply to diffusion models, see, e.g., [113] and [95, Chapter. 10] for an overview of the connection between the methods.
Consider a trainable velocity field $v(x,t;\theta)$ with parameters $\theta$, space (data) coordinates $x$ and time $t\in[0,1]$. The goal is to learn how samples from the source distribution, $\pi_{s}$, flow to the target distribution $\pi_{x}$. The source distribution is often chosen as a simple Gaussian, but can also be the distribution of observed data or corrupted images– which are of interest in the inverse problem context. The target distribution represents the QoI: physical model parameters or true images, for instance.
Before illustrating how to use FM for inverse problems, consider the specific training objective [96, 134] (various other forms exist)
$\min_{\theta}\mathbb{E}_{t\sim U[0,1],x\sim\pi_{x},y\sim\pi_{y}}\|v(x,t;\theta)-(x-y)\|_{2}^{2},$ (12)
where time $t$ is sampled uniformly $[0,1]$, and the $x$ and $y$ are sampled independently from each other, i.e, and unpaired and unsupervised setting. Integrating the velocity field in time (via the ODE $dx/dt=v(x,t;\theta)$ using e.g., forward Euler), starting with a sample $y\sim\pi_{y}$ will now generate a sample $x\sim\pi_{x}$. However, this does not mean that the generated sample from the QoI distribution will actually correspond to the sampled data $y$.
Several extensions to the basic FM formulation in Equation 12 are more suitable to solve inverse problems. If paired samples $(x,y)\sim\pi$ are available, we can train the velocity via
$\min_{\theta}\mathbb{E}_{t\sim U[0,1],(x,y)\sim\pi}\|v(x,t;\theta)-(x-y)\|_{2}^{2},$ (13)
which leads to promising results for image processing tasks like inpainting , super-resolution, and deblurring [4].
An extension to both Equation 12 and Equation 13 introduces a condition $c$, also referred to as a guidance signal [69, 120]. This condition may represent observed data, auxiliary information like an inpainting or data-acquisition mask, an image class-label, or even a textual description or caption. With additional conditions, Equation 13 becomes
$\min_{\theta}\mathbb{E}_{t\sim U[0,1],(x,y)\sim\pi(x,y|c)}\|v(x,t,c;\theta)-(x-y)\|_{2}^{2},$ (14)
which is similar to the formulation of [4], and where the paired data are sampled as $(x,y)\sim\pi(x,y|c)$, and the velocity field, $v(x,t,c;\theta)$, likely parameterized by a neural network, depends on $c$ as well. This technique also applies when the source distribution is chosen as Gaussian.
The above approaches require that the observed data $y$ and/or condition $c$ be available at the training time. When this is not the case, we may use techniques that include $y$ or $c$ at the inference stage. One such technique augments the velocity field with a guidance field $g(x,t,c)$ such that integrating the augmented field $v(x,t;\theta)+\alpha g(x,t,c)$ yields samples from the target distribution with desired properties, image-class, or adherence to observed data [125, 32, 113, 42]. Because $g(x,t,c)$ may be deterministic, or learned separately, the terms require balancing via the scalar $\alpha>0$.
@@
Assuming we have a forward differential operator, boundary conditions and spatial measurements $u(\chi_{j})=\tau_{j}$ for $j=1,\dots,m$, the solution of the PDE $u$ is modeled by a DNN, $\Phi(\,\cdot\,;\theta)$, parameterized by $\theta$. In the case of parameter estimation, these $\tau_{j}$ may coincide with our data $y_{i}$. Training simultaneously learns an approximation for the solution of the PDE $u(\chi)$ by $\Phi(\chi;\theta)$, and our QoI through minimization of the following quantities (if applicable): first, the mismatch computation of the forward model at the estimated solution; second, the violation of the predicted solution of boundary conditions; third the violation of the predicted solution of the physics governing the system; lastly, the mismatch between $u$ and $\tau_{j}$ evaluated at each $\chi_{j}$. Derivatives of $\Phi(\chi;\theta)$ can be computed using automatic differentiation, and mean-squared-error (MSE) is a common choice to compute these metrics [31].
@@
Method
$F$ for
training
$F$ for
inference
Data $y$
for training
Paired $\{x,y\}$
for training
@@
Flow - paired/conditional sampling
Flow - controlled generation
@@
Model-constrained approaches are similar to PINNs, but different in that they parametrize the inverse map by a DNN. Heuristically, these approaches learn the inverse mapping constrained by the forward map. Like PINNs, Model-constrained approaches attempt to ameliorate DNN data dependence by introducing additional information through regularization terms in optimization. Assuming a forward map $F$ is available, the model is aware that the observations it sees come from the forward model. Letting $\Phi(y;\theta)$ denote the learned inversion, model training optimizes
$\min_{\theta}\mathbb{E}_{(x,y)\sim\pi}\left\|x-\Phi(y;\theta)\right\|_{2}^{2}+\alpha\left\|y-F(\Phi(y;\theta))\right\|_{2}^{2}$ (15)
@@
Several strategies exist to represent $\Phi$. For example, [51] introduces the use of variational autoencoders and explores approximations of posterior distributions using this methodology. Further data-efficient approaches have been proposed in which the inverse map learns the solution to the Tikhonov-regularized inverse problem [108, 109].
@@ 1.3 Autoencoders
[…]
@@
The training of autoencoders is considered an unsupervised or self-supervised learning approach, where the network training learns to reconstruct the input data without relying on corresponding pre-defined labels or target data. The autoencoder’s learning process involves minimizing a loss function between the original input and the reconstructed output. For instance, for the mean squared error (MSE) loss, we have
$\min_{\theta}\ {\mathbb{E}_{x\sim\pi_{x}}\left\|d(e(x;\theta^{e});\theta^{d})-x\right\|_{2}^{2}.}$
@@ 1.4 Linear Autoencoder
In the case of linear autoencoders where both the encoder and decoder are linear transformations, i.e., $e=E\in\mathbb{R}^{r\times n}$ and $d=D\in\mathbb{R}^{n\times r}$ and each element of those matrices are trainable, with $r\leq n$ it is well known that the optimal solution is closely related to principal component analysis (PCA). This relationship has been established by assuming samples $x_{1},\ldots,x_{n_{x}}$ drawn independently from the distribution $\pi_{x}$ with $n_{x}\gg r$ and the corresponding data matrix $x_{1},\ldots,x_{n_{x}}]$ has at least column rank $r$. The mathematical equivalence between PCA and linear autoencoders has been established in several works, notably [20, 11, 112]. These studies demonstrate that the weight matrices of the encoder and decoder correspond to the eigenvectors of the data covariance matrix associated with the largest eigenvalues. This insight has been leveraged in various applications, including dimensionality reduction, denoising, and feature extraction, where linear autoencoders provide an interpretable and computationally efficient framework for data representation. For further details, see [14], which explores regularized extensions of linear autoencoders, and [145], which discusses their use in model reduction.
A deeper understanding of optimal linear autoencoders $A=DE\in\mathbb{R}^{n\times n}$ may be gained by treating $x$ as a random variable [^4] To simplify notation, we will use lowercase $x$ to represent a random variable, relying on context to distinguish between random variables and samples., and analyzing the problem through the lens of the Bayes risk formulation, i.e.,
$\min_{\text{rank}(A)\leq r}\ {\mathbb{E}_{x\sim\pi_{x}}}\left\|Ax-x\right\|_{2}^{2}.$ (16)
Here, $\mathbb{E}_{x\sim\pi_{x}}$} denotes the expectation of a random variable $x$. Let us assume that $x$ has a finite second moment that is symmetric positive definite $\Gamma_{x}=L_{x}L_{x}^{\top}=\mathbb{E}\,xx^{\top}$ with a symmetric decomposition by a matrix $L_{x}\in\mathbb{R}^{n}$. In this case, the objective function in Equation 16 reduces to
${\mathbb{E}_{x\sim\pi_{x}}}\left\|Ax-x\right\|_{2}^{2}$ $=\text{tr}\left((A-I)L_{x}L_{x}^{\top}(A-I)^{\top}\right)=\left\|AL_{x}-L_{x}\right\|_{\text{F}}^{2},$